# compat=1.0.0

### v1

@@ Abstract
Next-generation experiments, such as the Deep Underground Neutrino Experiment and the European Spallation Source, are set to dramatically improve sensitivity to neutron-antineutron oscillation that is a direct probe of $\Delta B=2$ baryon number violation. The discovery of such a rare process would indicate physics beyond the Standard Model and could point to specific unified theories that allow observable $n-\overline{n}$ transitions. We accordingly examine $n-\overline{n}$ oscillations within a unified framework that accounts for charged fermion masses and generates viable neutrino masses via the seesaw mechanism. More specifically, we show that $n-\overline{n}$ oscillations can arise from two specific topologies within two distinct $SU(5)$ scenarios. One topology requires a presence of two color-sextet scalars in the Type II seesaw framework, whereas the other involves a scalar sextet and a color-octet fermion in the Type III seesaw framework. While the former topology can be realized in the SO(10)SO(10)/Pati-Salam frameworks, the latter finds a natural embedding in $SU(5)$, which constitutes one of the key novelties of our work. Remarkably enough, the same dynamics responsible for fermion masses also induces baryon number violation, thus linking $n-\overline{n}$ oscillations to the flavor structure of the theory. We show that upcoming searches of such $\Delta B=2$ processes can probe new colored states with masses up to $10^{11}$ GeV that is well beyond the reach of colliders. This positions $n-\overline{n}$ oscillations as a rare low-energy portal to grand unification and ultra-heavy new physics.
@@ 1 Introduction
The observation of baryon number $(B)$ violation would provide unambiguous evidence of physics beyond the Standard Model (SM). Among the most intriguing possibilities predicted by grand unified theories (GUTs) [1, 2, 3, 4, 5, 6] are $\Delta B=2$ processes, such as neutron–antineutron ($n-\overline{n}$) oscillations. Although no such transitions have been observed to date, stringent experimental bounds already exist. In particular, Super-Kamiokande (SK) sets a strong lower limit on the $n-\overline{n}$ oscillation time given by $\tau_{n-\overline{n}}^{\mathrm{SK}}\geq 4.7\times 10^{8}\,\mathrm{s}$ [7]. On the other hand, ILL experiment [8] provides $\tau_{n-\overline{n}}^{\mathrm{ILL}}\geq 8.6\times 10^{7}\,\mathrm{s}$ from free $n-\overline{n}$ transitions. Future experiments are expected to significantly enhance this sensitivity. For example, the Deep Underground Neutrino Experiment (DUNE) [9] is expected to probe lifetimes exceeding $\tau_{n-\overline{n}}^{\mathrm{DUNE}}\sim 7\times 10^{8}$ s, while the proposed NNBAR experiment [10] at the European Spallation Source (ESS) aims to reach an impressive $\tau_{n-\overline{n}}^{\mathrm{NNBAR}}\sim 3\times 10^{9}$ s. Such unprecedented sensitivity enables the exploration of baryon number violating physics at energy scales far beyond the reach of current or foreseeable collider experiments, firmly establishing $n-\overline{n}$ oscillation searches as a uniquely powerful probe of high-scale new physics.
[…]
@@
Motivated by these remarkable developments and the growing experimental reach, in this work, we investigate two distinct topologies — Topology A and Topology B — which emerge naturally within unified theories, as shown in Fig. 1. In particular, we consider two simple models based on $SU(5)$ gauge group that generate viable masses for the SM charged fermions and neutrinos and simultaneously give rise to observable $n-\overline{n}$ oscillation signals. We demonstrate that the topologies shown in Fig.1 arise naturally by embedding the seesaw mechanism [11, 12, 13, 14, 15, 16, 17] for the neutrino mass generation within a $SU(5)$ grand unified framework. In this work, we consider two concrete realizations of these topologies, namely Model A and Model B, where Model A realizes Type II seesaw within the $SU(5)$ unified framework [18, 19, 20, 21, 22, 23, 24, 25], while Model B is based on a Type III seesaw embedding [26, 27, 28, 29]. Even though both Type II and Type III seesaws have been studied previously in the context of $SU(5)$ GUT, $n-\overline{n}$ oscillations within this context have not been explored, and the parameter space relevant for $n-\overline{n}$ oscillations has not been singled out. Our goal in this work is to rectify this omission. As we will show, when a Type II seesaw is embedded in a $SU(5)$ GUT, Topology A  [^1] Previously, Topology A of $n-\overline{n}$ oscillation has been studied primarily in the context of Pati–Salam and $SO(10)$ frameworks, or within SM extensions involving scalar diquarks [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]., which utilizes two distinct scalar sextets, naturally emerges. In contrast, a realization of a Type III seesaw gives rise to Topology B  [^2] Topology B has previously been explored in the context of SM extensions involving color-triplet scalars and right-handed neutrinos, as well as within certain supersymmetric frameworks [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]., involving a scalar sextet and a fermionic color octet. This particular possibility, to our knowledge, has not been considered in the literature before. Notably, the former scenario can be realized within the SO(10)SO(10)/Pati–Salam framework, while the latter finds a natural embedding in the $SU(5)$ setup, representing one of the key novelties of our work.
@@
The scenarios we consider are realistic since they successfully generate viable fermion masses, including neutrinos, achieve consistent gauge coupling unification, and satisfy current proton decay bounds, as well as present LHC and flavor physics constraints. Remarkably enough, the same dynamics responsible for the fermion mass generation also induces baryon-number violation, thereby linking $n-\overline{n}$ oscillations directly to the flavor structure of the theory. We demonstrate that upcoming searches for such $\Delta B=2$ processes can probe new colored states with masses up to $10^{11}$ GeV that is well beyond the reach of current or near-future colliders. This positions $n-\overline{n}$ oscillations as a unique low-energy window into grand unification and ultra-heavy new physics. We hope that the analysis presented in this work further motivates and inspires future experimental searches for $n-\overline{n}$ oscillations.
@@ 2.1 Charged fermion masses
The scalar sector that accomplishes the breaking of the $SU(3)\times SU(2)\times U(1)$ symmetry down to $SU(3)\times U(1)_{\mathrm{em}}$ and, consequentially, generates viable masses of the SM charged fermions and gauge bosons comprises a single 5-dimensional and a single 45-dimensional representation, where the properties of $45_{H}$, in the $SU(5)$ space, are $45^{\alpha\beta}_{H\gamma}=-45^{\beta\alpha}_{H\gamma}$ and $45^{\alpha\beta}_{H\alpha}=0$.
@@
such that $v_{5}^{2}+v_{45}^{2}=(246\,\mathrm{GeV})^{2}$. Here, we assume, without the loss of generality, that all the VEVs are real.
@@
where we explicitly specify contractions in both the $SU(5)$ and flavor spaces. These contractions, in turn, yield mass matrices at the GUT scale for the charged leptons, down-type quarks, and up-type quarks that are
@@
To generate viable $M_{N}$, we resort to two distinct models that, consequently, produce differing parametric dependences of the $n-\overline{n}$ oscillation signal. The first one uses a 15-dimensional scalar representation to produce neutrino masses via type II seesaw mechanism. We refer to it as Model A in what follows. The second model uses a 24-dimensional fermionic representation to generate neutrino masses via combination of type I and type III seesaw mechanisms. We refer to it as Model B. Note that both models generate charged fermion masses through the VEVs of $5_{H}$ and $45_{H}$, where the explicit forms of $M_{E}$, $M_{D}$, and $M_{U}$ are given in Eqs. (2.9), (2.10), and (2.11), respectively.
@@ 3.2 $n-\overline{n}$  oscillation
where $\widetilde{\mu}$ is a dimensionful parameter. We assume that $\widetilde{\mu}\lesssim 3\times\mathrm{max}(m_{\Delta_{6}},m_{\Sigma_{6}})$ to avoid the charge breaking minima. This issue, for example, has been studied in Refs. [67, 68, 69] for several specific scenarios.
@@
The $d=9$ operator that is relevant for $n-\overline{n}$ oscillations, in the $SU(3)\times U(1)_{\mathrm{em}}$ invariant basis, is given by [46]
@@
Following Ref. [46], we find the corresponding relevant hadronic matrix element for this operator to be $\langle\overline{n}|\mathcal{O}^{2}_{RRR}|n\rangle=-3\mathcal{M}_{1}(\mu)/20$, and take $\mathcal{M}_{1}(2\,\mathrm{GeV})=-4.6\times 10^{-4}\,\mathrm{GeV}^{6}$.
@@
We note that the current experimental bound, as given by Super-Kamiokande (SK) [7], is
@@
whereas the future DUNE [9] and NNBAR [10] experiments are expected to yield
@@
The lower bound in Eq. (3.20), as indicated, is given by the current SK limit [7], whereas the upper bound corresponds to the NNBAR expected sensitivity.
What we need to address next is whether Model A can indeed yield this window of observability.
@@ 3.3.2 Constraints from collider data and flavor physics
We note that the LHC mass limits on scalars that couple to quark-quark pairs are rather stringent and can go up to $7.5$ TeV [70] depending on the associated branching ratios. We take these limits into account when we perform our numerical fit and check for the Model A viability. Also, in Model A we have scalar color sextets $\Delta_{6}$ and $\Sigma_{6}$ that couple to $d^{c}$-$d^{c}$ and $u^{c}$-$d^{c}$ pairs, respectively. This, consequentially, can lead to meson-antimeson oscillations through diagrams given in Figs. 4 and 5. For the $\Delta_{6}$ sextet, tree-level processes provide the most stringent constraints on the parameter space, while box diagrams put bounds on different combinations of Yukawa couplings. In contrast, for the $\Sigma_{6}$ sextet, no such tree-level processes are relevant and box diagrams provide the leading bounds. We implement constraints from these processes in our numerical analysis. The limits in question are taken from Ref. [71] (for model independent bounds, see, for example, Ref. [72]) and summarized in Table 1.
[…]
@@
Table 1: Relevant flavor physics bounds inferred from Ref. [71] (for model independent bounds, see, for example, Ref. [72]). For the $\Delta_{6}$ sextet, tree-level processes provide the most stringent constraints on the parameter space, while box diagrams put bounds on different combinations of Yukawa couplings. In contrast, for the $\Sigma_{6}$ sextet, no such tree-level processes are relevant and box diagrams provide the leading constraints.
@@ 3.3.3 Numerical Fit
Fermion mass fit, for Model A, is performed at the gauge coupling unification scale $M_{\mathrm{GUT}}$, with the input values at $M_{\mathrm{GUT}}$ that are taken from Ref. [73] for convenience. We list these input values in Table 2, as well as the associated outcome of our numerical fit. Note that we use a freedom to freely rotate $10_{Fi}$ and $\overline{5}_{Fj}$ representations in the $SU(5)$ space to go into basis where $Y_{1}$ of Eqs. (2.12) and (2.13) is a diagonal matrix with real parameters. We perform a simultaneous fit of the charged lepton, down-type quark, and neutrino sectors. Whereas, in the up-type quark sector, due to $M_{U}=M_{U}^{T}$, the unitary matrices $U_{c}$ and $U$ are completely determined via Eq. (3.21). Of course, what interests us the most are the explicit values of $Y_{11}^{dd}$ and $Y_{11}^{ud}$ as these are crucial input parameters for the strength of the $n-\overline{n}$ oscillation signal.
@@
The outcome of our numerical fit for the Model A (IO) scenario is specified in a fourth column of Table 2. Note that this scenario can be fully tested by near future neutrino observables. Namely, neutrino masses are predicted to be $\left(m_{1},m_{2},m_{3}\right)=\left(49.64,50.39,5.36\right)\,\mathrm{meV}$, which leads to somewhat large value for $\sum_{i}m_{i}$, which is close to the ruled out bound from cosmological data [74] that suggests $\sum m_{i}<87$ meV or $\sum m_{i}<120$ meV, depending on experiments included. Model A (IO) fit furthermore yields $m_{\beta\beta}=48.02$ meV, which will also be probed in very near future. Furthermore, it is exciting to point out that even though we do not attempt to fit the CP violating phase in the leptonic sector, yet, our benchmark fit corresponds to $\delta^{\mathrm{PMNS}}_{\mathrm{CP}}=290^{\circ}$ which is well within the $1\,\sigma$ range of the current experimental global fit of all neutrino observables.
[…]
@@
$\sum_{i}m_{i}$ (meV) <(8787120120[74] 75.08 105.39
$m_{\beta\beta}$ (meV) <(2828122122[75] 10.13 48.02
$\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$ (deg) $177^{+19}_{-20}$$285^{+25}_{-28}$ [76] 74.7 290.489
Table 2: Charged fermion mass parameter input at $M_{\mathrm{GUT}}$ is taken from Ref.  [73]. Neutrino observables are taken from Refs. [77, 76]. The Model A (NO) fit yields $\chi^{2}=1.53$, whereas Model A (IO) fit gives $\chi^{2}=0.7$. Some of the interesting fit outcomes relevant for current and future neutrino experiments are also summarized in this Table. These quantities are the sum of neutrino masses $\sum_{i}m_{i}$, the neutrinoless double beta decay parameter $m_{\beta\beta}=|\sum_{i}(U_{\mathrm{PMNS}})^{2}_{ei}m_{i}|$, and the CP-violating Dirac phase in the leptonic sector $\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$.
@@
Figure 6: Observable $n-\overline{n}$ range for the Model A (NO) and Model A (IO) scenarios, as indicated. The shaded blue and magenta regions correspond to the viable parts of the parameter space that can be probed in the future experiments, namely, through NNBAR and DUNE experiments, respectively, whereas the shaded gray region plotted on top of the magenta and blue regions shows the current SK exclusion. Light-green vertical shaded region shows the part excluded by $K^{0}-\overline{K}^{0}$ oscillation. Moreover, the pink shaded part highlights the current LHC bound due to dijet resonance. NO/IO here refers to normal/inverted mass ordering for neutrinos. Both panels are obtained for $\tilde{\mu}=m_{\Delta_{6}}$.
We have already mentioned that the phenomenologically viable gauge coupling unification requires $\Sigma_{5}(3,3,-1/3)\in 45_{H}$ to be significantly below the unification scale if both $\Sigma_{6}$ and $\Delta_{6}$ scalars are light. As the mass ranges of both $\Sigma_{6}$ and $\Delta_{6}$ are explicitly given in Fig. 6, we can finally quantify this statement by performing the one-loop level gauge coupling unification analysis with the following SM input parameters: $M_{Z}=91.1876$ GeV, $\alpha_{S}(M_{Z})=0.1193\pm 0.0016$, $\alpha^{-1}(M_{Z})=127.906\pm 0.019$, and $\sin^{2}\theta_{W}(M_{Z})=0.23126\pm 0.00005$ [78].

### v2

# Beyond Neutrino Mass: Observable $𝒏¯\overline{n}$ Oscillations in UV Complete Seesaw Models

@@ Abstract
Next-generation experiments, such as the Deep Underground Neutrino Experiment and the European Spallation Source, are set to improve sensitivity to neutron-antineutron oscillation, a direct probe of $\Delta B=2$ baryon number violation, with particularly significant gains expected at the latter. The discovery of such a rare $\Delta B=2$ process would indicate physics beyond the Standard Model and could point to specific unified theories that allow observable $n-\overline{n}$ transitions. We accordingly examine $n-\overline{n}$ oscillations within a unified framework that accounts for charged fermion masses and generates viable neutrino masses via the seesaw mechanism. More specifically, we show that $n-\overline{n}$ oscillations can arise from two specific topologies within two distinct $SU(5)$ scenarios. One topology requires a presence of two color-sextet scalars in the Type II seesaw framework, whereas the other involves a scalar sextet and a color-octet fermion in the Type III seesaw framework. While the former topology can be realized in the SO(10)SO(10)/Pati-Salam frameworks, the latter finds a natural embedding in $SU(5)$, which constitutes one of the key novelties of our work. Remarkably enough, the same dynamics responsible for fermion masses also induces baryon number violation, thus linking $n-\overline{n}$ oscillations to the flavor structure of the theory. We show that, given a TeV-scale mass for one of the colored states, upcoming searches for such $\Delta B=2$ processes can probe for a presence of the other colored states with masses up to $10^{11}\,\mathrm{GeV}$, well beyond the reach of colliders. This positions $n-\overline{n}$ oscillations as a rare low-energy portal to grand unification and ultra-heavy new physics.
@@ 1 Introduction
The observation of baryon number $(B)$ violation would provide unambiguous evidence of physics beyond the Standard Model (SM). Among the most intriguing consequences of grand unified theories (GUTs) Pati:1973rp , Pati:1974yy , Georgi:1974sy , Georgi:1974yf , Georgi:1974my , Fritzsch:1974nn are baryon-number-violating processes. While the hallmark prediction of GUTs is $\Delta B=1$ proton decay, for example $p\to e^{+}\pi^{0}$, $\Delta B=2$ processes, such as neutron–antineutron ($n-\overline{n}$) oscillations, may dominate in certain models or specific regions of the parameter space. Although no such $n-\overline{n}$ transitions have been observed to date, stringent experimental bounds already exist. In particular, Super-Kamiokande (SK) sets a strong lower limit on the $n-\overline{n}$ oscillation time given by $\tau_{n-\overline{n}}^{\mathrm{SK}}\geq 4.7\times 10^{8}\,\mathrm{s}$ Super-Kamiokande:2020bov . On the other hand, ILL experiment Baldo-Ceolin:1994hzw provides $\tau_{n-\overline{n}}^{\mathrm{ILL}}\geq 8.6\times 10^{7}\,\mathrm{s}$ from free $n-\overline{n}$ transitions. Future experiments are expected to significantly enhance this sensitivity. For example, the Deep Underground Neutrino Experiment (DUNE) DUNE:2015lol is expected to probe lifetimes exceeding $\tau_{n-\overline{n}}^{\mathrm{DUNE}}\sim 7\times 10^{8}$ s, while the proposed NNBAR experiment Addazi:2020nlz at the European Spallation Source (ESS) aims to reach an impressive $\tau_{n-\overline{n}}^{\mathrm{NNBAR}}\sim 3\times 10^{9}$ s. Such unprecedented sensitivity enables the exploration of baryon number violating physics at energy scales far beyond the reach of current or foreseeable collider experiments, firmly establishing $n-\overline{n}$ oscillation searches as a uniquely powerful probe of high-scale new physics.
$S_{2}$$S_{1}$$S_{3}$$u^{c}$$d^{c}$$u^{c}$$d^{c}$$d^{c}$$d^{c}$
@@
Motivated by these remarkable developments and the growing experimental reach, in this work, we investigate two distinct topologies — Topology A and Topology B — which emerge naturally within unified theories, as shown in Fig. 1. In particular, we consider two simple models based on $SU(5)$ gauge group that generate viable masses for the SM charged fermions and neutrinos and simultaneously give rise to observable $n-\overline{n}$ oscillation signals. We demonstrate that the topologies shown in Fig.1 arise naturally by embedding the seesaw mechanism Minkowski:1977sc , Yanagida:1979as , Glashow:1979nm , Gell-Mann:1979vob , Mohapatra:1979ia , Schechter:1980gr , Schechter:1981cv for the neutrino mass generation within an $SU(5)$ grand unified framework. In this work, we consider two concrete realizations of these topologies, namely Model A and Model B, where Model A realizes Type II seesaw within the $SU(5)$ unified framework Ma:1998dn , Dorsner:2005fq , Dorsner:2005ii , Dorsner:2006hw , Dorsner:2007fy , Antusch:2022afk , Calibbi:2022wko , Antusch:2023mqe , Kaladharan:2024bop , Dev:2025sox , while Model B is based on a Type III seesaw embedding Ma:1998dn , Bajc:2006ia , Dorsner:2006fx , FileviezPerez:2007bcw , Antusch:2023kli . Even though both Type II and Type III seesaws [^1] Within the $SU(5)$ framework, the type I seesaw mechanism is less appealing, as it requires introduction of gauge-singlet states. have been studied previously in the context of $SU(5)$ GUT, $n-\overline{n}$ oscillations within this context have not been explored, and the parameter space relevant for $n-\overline{n}$ oscillations has not been singled out. Our goal in this work is to rectify this omission. As we will show, when a Type II seesaw is embedded in a $SU(5)$ GUT, Topology A  [^2] Previously, Topology A of $n-\overline{n}$ oscillation has been studied primarily in the context of Pati–Salam and $SO(10)$ frameworks, or within SM extensions involving scalar diquarks Mohapatra:1980qe , Mohapatra:1982xz , Chang:1984qr , Babu:2006xc , Babu:2008rq , Baldes:2011mh , Babu:2012vc , Babu:2012qia , Arnold:2012sd , Bernal:2012gv , Babu:2013yca , Patra:2014goa , Herrmann:2014fha , Saad:2017pqj , Asaka:2019ocw , Girmohanta:2020qfd , Fridell:2021gag , Patel:2022nek ., which utilizes two distinct scalar sextets, naturally emerges. In contrast, a realization of a Type III seesaw gives rise to Topology B  [^3] Topology B has previously been explored in the context of SM extensions involving color-triplet scalars and right-handed neutrinos, as well as within certain supersymmetric frameworks Zwirner:1983dgv , Barbieri:1985ty , Mohapatra:1986bd , Lazarides:1986jt , Goity:1994dq , Babu:2001qr , Babu:2006wz , Allahverdi:2010im , Gu:2011ff , Gu:2011fp , Allahverdi:2013mza , Dev:2015uca , Dhuria:2015swa , Ghalsasi:2015mxa , Gu:2016ghu , Calibbi:2016ukt , Gu:2017cgp , Calibbi:2017rab , Allahverdi:2017edd , Grojean:2018fus ., involving a scalar sextet and a fermionic color octet. This particular possibility, to our knowledge, has not been considered in the literature before. Notably, the former scenario can be realized within the SO(10)SO(10)/Pati–Salam framework, while the latter finds a natural embedding in the $SU(5)$ setup, representing one of the key novelties of our work.
@@
The scenarios we consider are realistic since they successfully generate viable fermion masses, including neutrinos, achieve consistent gauge coupling unification, and satisfy current proton decay bounds, as well as present LHC and flavor physics constraints. Remarkably enough, the same dynamics responsible for the fermion mass generation also induces baryon-number violation, thereby linking $n-\overline{n}$ oscillations directly to the flavor structure of the theory. We demonstrate that, given a TeV-scale mass for one of the colored states, upcoming searches for such $\Delta B=2$ processes can probe the other colored states with masses up to $10^{11}\,\mathrm{GeV}$, well beyond the reach of current or near-future colliders. This positions $n-\overline{n}$ oscillations as a unique low-energy window into grand unification and ultra-heavy new physics. We hope that the analysis presented in this work further motivates and inspires future experimental searches for $n-\overline{n}$ oscillations.
@@ 2.1 Charged fermion masses
The scalar sector that accomplishes the breaking of the $SU(3)\times SU(2)\times U(1)$ symmetry down to $SU(3)\times U(1)_{\mathrm{em}}$ and, consequentially, generates viable masses of the SM charged fermions and gauge bosons comprises a single 5-dimensional and a single 45-dimensional representation [^4] An alternative option to correct the wrong mass relations is to add a pair of $5+\overline{5}$ Babu:2012pb or $15+\overline{15}$ Oshimo:2009ia , Dorsner:2019vgf vectorlike fermions. However, as will be shown later, the presence of a $45_{H}$ in our setup is essential not only to achieve gauge coupling unification but to also generate observable $n-\overline{n}$ oscillations. , where the properties of $45_{H}$, in the $SU(5)$ space, are $45^{\alpha\beta}_{H\gamma}=-45^{\beta\alpha}_{H\gamma}$ and $45^{\alpha\beta}_{H\alpha}=0$.
@@
such that $|v_{5}|^{2}+|v_{45}|^{2}=(246\,\mathrm{GeV})^{2}$. Without loss of generality, one can always take $v_{5}$ to be real. On the other hand, $v_{45}$ can still be complex. However, its phase is irrelevant for the fermion mass fit, as it can always be absorbed into the corresponding Yukawa coupling matrices. Therefore, for simplicity, we take $v_{45}$ also to be real.
@@
where we explicitly specify contractions in both the $SU(5)$ and flavor spaces ($\varepsilon$ here is the five-index Levi-Civita tensor). These contractions, in turn, yield mass matrices at the GUT scale for the charged leptons, down-type quarks, and up-type quarks that are
@@
To generate viable $M_{N}$, we resort to two distinct models that, consequently, produce differing parametric dependence of the $n-\overline{n}$ oscillation signal. The first one uses a 15-dimensional scalar representation to produce neutrino masses via type II seesaw mechanism. We refer to it as Model A in what follows. The second model uses a 24-dimensional fermionic representation to generate neutrino masses via combination of type I and type III seesaw mechanisms. We refer to it as Model B. Note that both models generate charged fermion masses through the VEVs of $5_{H}$ and $45_{H}$, where the explicit forms of $M_{E}$, $M_{D}$, and $M_{U}$ are given in Eqs. (2.9), (2.10), and (2.11), respectively.
@@ 3.2 $n-\overline{n}$  oscillation
where $\widetilde{\mu}$ is a dimensionful parameter. We assume that $\widetilde{\mu}\lesssim 3\times\mathrm{max}(m_{\Delta_{6}},m_{\Sigma_{6}})$ to avoid the charge breaking minima. This issue, for example, has been studied in Refs. Frere:1983ag , Alvarez-Gaume:1983drc , Casas:1996de for several specific scenarios.
@@
The $d=9$ operator that is relevant for $n-\overline{n}$ oscillations, in the $SU(3)\times U(1)_{\mathrm{em}}$ invariant basis, is given by Fridell:2021gag
@@
Following Ref. Fridell:2021gag , we find the corresponding relevant hadronic matrix element for this operator to be $\langle\overline{n}|\mathcal{O}^{2}_{RRR}|n\rangle=-3\mathcal{M}_{1}(\mu)/20$, and take $\mathcal{M}_{1}(2\,\mathrm{GeV})=-4.6\times 10^{-4}\,\mathrm{GeV}^{6}$.
@@
We note that the current experimental bound, as given by Super-Kamiokande (SK) Super-Kamiokande:2020bov , is
@@
whereas the future DUNE DUNE:2015lol and NNBAR Addazi:2020nlz experiments are expected to yield
@@
The lower bound in Eq. (3.20), as indicated, is given by the current SK limit Super-Kamiokande:2020bov , whereas the upper bound corresponds to the NNBAR expected sensitivity.
What we need to address next is whether Model A can indeed yield this window of observability, i.e., $4.7\times 10^{8}\,\mathrm{s}\leq\tau_{n-\overline{n}}\leq 3\times 10^{9}\,\mathrm{s}$.
@@ 3.3.2 Constraints from collider data and flavor physics
We note that the LHC mass limits on scalars that couple to quark-quark pairs are rather stringent and can go up to $7.5$ TeV CMS:2019gwf depending on the associated branching ratios. Interestingly, future projections for such diquark resonances at the High-Luminosity LHC (HL-LHC) and the Future Circular Collider (FCC-hh) extend the discovery prospect up to $9.4$ TeV and $63$ TeV for integrated luminosities of $3$ ab-1 and $30$ ab-1 Harris:2022kls , Bernardi:2022hny , respectively. We take these limits into account when we perform our numerical fit and check for the Model A viability. Also, in Model A we have scalar color sextets $\Delta_{6}$ and $\Sigma_{6}$ that couple to $d^{c}$-$d^{c}$ and $u^{c}$-$d^{c}$ pairs, respectively. This, consequentially, can lead to meson-antimeson oscillations through diagrams given in Figs. 4 and 5. For the $\Delta_{6}$ sextet, tree-level processes provide the most stringent constraints on the parameter space, while box diagrams put bounds on different combinations of Yukawa couplings. In contrast, for the $\Sigma_{6}$ sextet, no such tree-level processes are relevant and box diagrams provide the leading bounds. We implement constraints from these processes in our numerical analysis. The limits in question are taken from Ref. Fortes:2013dba (for model independent bounds, see, for example, Ref. Isidori:2023sma ) and summarized in Table 1.
$\Delta_{6}$$d^{c}$$d^{c}$$s^{c}$$s^{c}$$Y^{dd}_{11}$$Y^{dd}_{22}$$^{*}$
@@
Table 1: Relevant flavor physics bounds inferred from Ref. Fortes:2013dba (for model independent bounds, see, for example, Ref. Isidori:2023sma ). For the $\Delta_{6}$ sextet, tree-level processes provide the most stringent constraints on the parameter space, while box diagrams put bounds on different combinations of Yukawa couplings. In contrast, for the $\Sigma_{6}$ sextet, no such tree-level processes are relevant and box diagrams provide the leading constraints.
@@ 3.3.3 Numerical Fit
Fermion mass fit, for Model A, is performed at the gauge coupling unification scale $M_{\mathrm{GUT}}$, with the input values at $M_{\mathrm{GUT}}$ that are taken from Ref. Babu:2016bmy for convenience. We list these input values in Table 2, as well as the associated outcome of our numerical fit. Note that we use a freedom to freely rotate $10_{Fi}$ and $\overline{5}_{Fj}$ representations in the $SU(5)$ space to go into basis where $Y_{1}$ of Eqs. (2.12) and (2.13) is a diagonal matrix with real parameters. We perform a simultaneous fit of the charged lepton, down-type quark, and neutrino sectors. Whereas, in the up-type quark sector, due to $M_{U}=M_{U}^{T}$, the unitary matrices $U_{c}$ and $U$ are completely determined via Eq. (3.21). Of course, what interests us the most are the explicit values of $Y_{11}^{dd}$ and $Y_{11}^{ud}$ as these are crucial input parameters for the strength of the $n-\overline{n}$ oscillation signal.
@@
The outcome of our numerical fit for the Model A (IO) scenario is specified in a fourth column of Table 2. Note that this scenario can be fully tested by near future neutrino observables. Namely, neutrino masses are predicted to be $\left(m_{1},m_{2},m_{3}\right)=\left(49.64,50.39,5.36\right)\,\mathrm{meV}$, which leads to somewhat large value for $\sum_{i}m_{i}$, which is close to the ruled out bound from cosmological data ParticleDataGroup:2024cfk that suggests $\sum m_{i}<87$ meV or $\sum m_{i}<120$ meV, depending on experiments included. Model A (IO) fit furthermore yields $m_{\beta\beta}=48.02$ meV, which will also be probed in very near future. Furthermore, it is exciting to point out that even though we do not attempt to fit the CP violating phase in the leptonic sector, yet, our benchmark fit corresponds to $\delta^{\mathrm{PMNS}}_{\mathrm{CP}}=290^{\circ}$ which is well within the $1\,\sigma$ range of the current experimental global fit of all neutrino observables.
Model A Experimental values (NO/IO) Fit values (NO) Fit values (IO)
@@
$\sum_{i}m_{i}$ (meV) <(8787120120ParticleDataGroup:2024cfk 75.08 105.39
$m_{\beta\beta}$ (meV) <(2828122122KamLAND-Zen:2024eml 10.13 48.02
$\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$ (deg) $177^{+19}_{-20}$$285^{+25}_{-28}$ NUFIT 74.7 290.489
Table 2: Charged fermion mass parameter input at $M_{\mathrm{GUT}}$ is taken from Ref. Babu:2016bmy . Neutrino observables are taken from Refs. Esteban:2024eli , NUFIT . The Model A (NO) fit yields $\chi^{2}=1.53$, whereas Model A (IO) fit gives $\chi^{2}=0.7$. Some of the interesting fit outcomes relevant for current and future neutrino experiments are also summarized in this Table. These quantities are the sum of neutrino masses $\sum_{i}m_{i}$, the neutrinoless double beta decay parameter $m_{\beta\beta}=|\sum_{i}(U_{\mathrm{PMNS}})^{2}_{ei}m_{i}|$, and the CP-violating Dirac phase in the leptonic sector $\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$.
@@
Figure 6: Observable $n-\overline{n}$ range for the Model A (NO) and Model A (IO) scenarios, as indicated. The shaded blue and magenta regions correspond to the viable parts of the parameter space that can be probed in the future experiments, namely, through NNBAR and DUNE experiments, respectively, whereas the shaded gray region plotted on top of the magenta and blue regions shows the current SK exclusion. Light-green vertical shaded region shows the part excluded by $K^{0}-\overline{K}^{0}$ oscillation. Moreover, the pink shaded part highlights the current LHC bound due to dijet resonance. Projected sensitivity for dijet resonances from FCC-$h$ is presented by the dash-dotted red line, which will probe a significant part of the existing parameter space. NO/IO here refers to normal/inverted mass ordering for neutrinos. Both panels are obtained for $\tilde{\mu}=m_{\Delta_{6}}$.
We have already mentioned that the phenomenologically viable gauge coupling unification requires $\Sigma_{5}(3,3,-1/3)\in 45_{H}$ to be significantly below the unification scale if both $\Sigma_{6}$ and $\Delta_{6}$ scalars are light. As the mass ranges of both $\Sigma_{6}$ and $\Delta_{6}$ are explicitly given in Fig. 6, we can finally quantify this statement by performing the one-loop level gauge coupling unification analysis with the following SM input parameters: $M_{Z}=91.1876$ GeV, $\alpha_{S}(M_{Z})=0.1193\pm 0.0016$, $\alpha^{-1}(M_{Z})=127.906\pm 0.019$, and $\sin^{2}\theta_{W}(M_{Z})=0.23126\pm 0.00005$ Agashe:2014kda .